Peninsula Technologies Co. is a U.S. firm that sells electronic equipment. One of its foreign subsidiaries had $3.4 m of taxable income in the fiscal year. Calculate what the foreign and U.S. taxes would be on this income in the two cases presented below:

Baseline Values : Case 1 /Case 2

a. Foreign corporate income tax rate 28% / 40%

b. U.S. corporate income tax rate 35% / 35%

c. Foreign dividend withholding tax rate 0% / 10%

d. U.S. ownership in foreign firm 100% / 100%

e. Dividend payout rate of foreign firm 100% / 100%

Q1. What are the total taxes paid and the effective tax rate in both cases?

Q2. If the foreign income is not repatriated to the U.S., what would be the total taxes paid and

the effective tax rate?
Answer #1

Case - 1: If the foreign income is repatriated to the US:

a. Foreign corporate income tax rate - 28%

b. U.S. corporate income tax rate - 35%

c. Foreign dividend withholding tax rate - 0%

Calculation of foreign taxes:

Taxable income of a foreign subsidiary - $3.4 m

Less: Foreign corporate income tax rate @28% - $0.952 m

Income declared as dividend - $2.448 m

Less: Foreign dividend withholding tax rate @ 0% - $0.000 m

Income available for shareholders         - $2.448 m

Calculation of U.S. taxes:

Foreign income part of US Income   - $3.4 m

U.S. corporate income tax rate @35%   - $1.19 m

Less: Foreign corporate income tax already paid - $ 0.952 m

Less: Foreign dividend withholding tax already paid - $ 0.00 m

Additional US taxes due - $ 0.238 m

Total taxes to be paid ($0.952 m + $0.238 m) - $1.19 m

Income remaining in US - $2.21 m

Effective tax-rate on Foreign Income - (0.952/3.4)*100 = 28%

Effective tax-rate on US Income - (1.19/3.4)*100 = 35%

Case - 2: If the foreign income is repatriated to the US:

a. Foreign corporate income tax rate - 40%

b. U.S. corporate income tax rate - 35%

c. Foreign dividend withholding tax rate - 10%

Calculation of foreign taxes:

Taxable income of a foreign subsidiary - $3.400 m

Less: Foreign corporate income tax rate @40% - $1.360 m

Income declared as dividend - $2.040 m

Less: Foreign dividend withholding tax rate @10%        - $0.204 m

Income available for shareholders     - $1.836 m

Calculation of U.S. taxes:

Foreign income part of US Income   - $3.400 m

U.S. corporate income tax rate @35%   - $1.190 m

Less: Foreign corporate income tax already paid - $1.360 m

Less: Foreign dividend withholding tax already paid - $0.204 m

Additional US taxes due - None

Total taxes to be paid ($1.360 m + $0.204 m) - $1.564 m

Income remaining in US - $1.836 m

Effective tax-rate on Foreign Income - (1.564/3.4)*100 = 46%

Effective tax-rate on US Income - (1.564/3.4)*100 = 46%

Case - 1: If the foreign income is not repatriated to the US:

Total taxes paid - $0.952 m

Effective tax-rate on Foreign Income - (0.952/3.4)*100 = 28%

Case - 2: If the foreign income is not repatriated to the US:

Total taxes paid - $0.952 m

Effective tax-rate on Foreign Income - (1.364/3.4)*100 = 40%